Let 𝑋~𝑁(40,144) and if π‘Œ = 2𝑋 βˆ’ 1

Find the following probabilities:

(i) 𝑃(𝑋 ≀ 40)

(ii) 𝑃(π‘Œ β‰₯ 50)

(iii)𝑃(35 ≀ 𝑋 ≀ 45)

(iv)𝑃(βˆ’45 ≀ π‘Œ ≀ 45)

(v)𝑃(βˆ’50 ≀ π‘Œ ≀ 100)


1
Expert's answer
2021-04-14T07:38:49-0400

If 𝑋~𝑁(40,144) then the mathematical expectation is "a(X)=40", variance is "{\sigma ^2}(X) = 144 \Rightarrow \sigma (X) = 12" .

Then

"a(Y) = a(2X - 1) = 2a(X) - a(1) = 2 \cdot 40 - 1 = 79"

"{\sigma ^2}(Y) = {\sigma ^2}(2X - 1) = {2^2}{\sigma ^2}(X) - {\sigma ^2}(1) = 4{\sigma ^2}(X) = 4 \cdot 144 = {\rm{576}} \Rightarrow \sigma (Y) = 24"

(i) Let's find

"P(X \le 40) = \Phi \left( {\frac{{\beta - a(X)}}{{\sigma (X)}}} \right) - \Phi \left( { - \infty } \right) = \Phi \left( {\frac{{40 - 40}}{{12}}} \right) + 0.5 = 0 + 0.5 = 0.5"

Answer: "P(X \le 40) = 0.5"

(ii) "P(Y \ge 50) = \Phi \left( \infty \right) - \Phi \left( {\frac{{\alpha - a(Y)}}{{\sigma (Y)}}} \right) = 0.5 - \Phi \left( {\frac{{50 - 79}}{{24}}} \right) = 0.5 + \Phi \left( {1.2} \right) = 0.5 + 0.3849 = 0.8849"

Answer: "P(Y \ge 50) = 0.8849"

(iii) "P(35 \le X \le 45) = \Phi \left( {\frac{{\beta - a(X)}}{{\sigma (X)}}} \right) - \Phi \left( {\frac{{\alpha - a(X)}}{{\sigma (X)}}} \right) = \Phi \left( {\frac{{45 - 40}}{{12}}} \right) + \Phi \left( {\frac{{35 - 40}}{{12}}} \right) = \Phi (0.42) + \Phi (0.42) = 2 \cdot 0.1628 = 0.3256"

Answer: "P(35 \le X \le 45) = 0.3256"

(iv) "P( - 45 \le Y \le 45) = \Phi \left( {\frac{{\beta - a(Y)}}{{\sigma (Y)}}} \right) - \Phi \left( {\frac{{\alpha - a(Y)}}{{\sigma (Y)}}} \right) = \Phi \left( {\frac{{45 - 79}}{{24}}} \right) - \Phi \left( {\frac{{ - 45 - 79}}{{24}}} \right) = - \Phi (1.42) + \Phi (5.17) = - 0.4222 + 0.5 = 0.0778"

Answer: "P( - 45 \le Y \le 45) = 0.0778"

(v) "P( - 50 \le Y \le 100) = \Phi \left( {\frac{{100 - 79}}{{24}}} \right) - \Phi \left( {\frac{{ - 50 - 79}}{{24}}} \right) = \Phi (0.875) + \Phi (5.375) = 0.3092 + 0.5 = 0.8092"

Answer: "P( - 50 \le Y \le 100) = 0.8092"